# Speeding-up Large-scale LP Energy System Models: Using Graph-theory to Remove the Overhead Cost of Flexible Modeling

### v2

@@ 3.2 Two building blocks with two flow variables (2BB-2F)
3.3 Two building blocks with one flow variable (2BB-1F)
This approach uses as building blocks the energy assets and the balance nodes. in addition, it uses a single free variable to represent the flow between nodes, which can take positive and negative values, instead of two positive variables as in the previous approach. The main advantage of this method is that it eliminates the possibility of bidirectional flow between two nodes, as there is only one variable. However, it is not possible to model transmission or distribution networks losses using a single free variable for the flow. Lastly, the free variable must have bounds in both directions to represent the capacity limits between the two nodes.
@@ 3.3 Two building blocks with one flow variable (2BB-1F)
3.4 One building block with one flow variable (1BB-1F)
This proposed approach only uses as building blocks the energy assets. Taking advantage of the graph-theory principles establishes the connection between energy assets as vertices and energy flows as edges. Connecting assets directly to each other (without any intervening nodes) can significantly reduce the number of variables and constraints required to represent the system. When transmission or distribution network losses are significant, two conversion assets can be utilised to represent them—one for each direction with input-output ratios.. A similar approach can be adopted for gas pressure flows. Therefore, this method enables us to easily model simple and complex situations, reducing the model size while retaining the same accuracy.
Figure 1(d) shows the battery with a renewable example for the proposed 1BB-1F. Since the assets can connect among them, the battery can directly charge from renewable, and there is no need for an extra constraint to avoid charging from the grid. Therefore, we need 4 variables (3 flow variables + 1 storage level) and 6 constraints (4 capacity limits + 1 demand balance + 1 storage balance).
@@ 3.4 One building block with one flow variable (1BB-1F)
: flow from the battery to the electricity demand balance with a maximum capacity
: flow from the solar pv to the battery with a maximum capacity
: flow from the solar pv to the electricity demand balance with a maximum capacity
3.5 Summary
Table 1 summarises the number of variables and constraints per time step for each modelling approach in the example shown in Figure 1. It also shows the reduction in the number of variables and constraints, with the 3BB-4F approach as a reference.
@@ 3.5 Summary
| Modelling Approach | Variables | Constraints | Nonzeros |
@@ 4.1 Energy system optimisation model
TulipaEnergyModel.jl is an optimisation model using 1BB-1F and determines the optimal investment and operation decisions for different types of assets (e.g., producers, consumers, conversion, storage, and transport). It is developed in Julia [16] and depends mainly on the JuMP.jl [17] and Graphs.jl [18] packages.
The complete description of the model, its core concepts, mathematical formulation, and tutorials are available in the GitHub documentation of the model 111https://tulipaenergy.github.io/TulipaEnergyModel.jl/stable/.
4.2 Case study
@@
Figure 2 shows an illustrative integrated energy system with three interconnected areas named Asgard, Midgard, and Valhalla. The diagram uses the 2BB-2F approach. In addition, it includes the flow and balance of electricity, heat, and gas within a mock-up energy grid to explore different possibilities for flows among energy assets using nodes. Asgard includes a combined cycle gas turbine, a solar photovoltaic installation, and a battery system. Midgard features a wind park, a hydro plant, and a small modular reactor for nuclear power generation. Valhalla focuses on hydrogen as an energy vector, with a hydrogen generator, a hydrogen storage facility, and a fuel cell. Transmission lines and gas pipelines connect the system, allowing energy transfer between areas.
Figures 3 and 4 show the equivalent energy system for the other two approaches. In general, each arrow represent a variable and each BB element represents a constraint. Note that the number of flows (arrows) reduces compared to the 2BB-2F approach. For the purpose of evaluating the impact of different problem sizes, we created six instances of the problem, ranging from small to large-scale optimisation, with the smallest being labelled as 1 and the largest as 6. Section 5 analyses the impact of these reductions in the model for each approach and instance. Finally, Section 9 has the link to the repository with the input data files for each approach.
@@ 5 Results
The results in this section were obtained using TulipaEnergyModel.jl version 0.6.1 and Gurobi version 11.0.0 on a 12th Gen Intel(R) Core(TM) i7-1255U 1.70 GHz processor, and 16.0 GB RAM. Section 9 includes the link to the repository with the files to reproduce the experiments in this paper.
Table 2 shows the objective functions and problem size for each instance, where 2BB-2F is used as a reference in the table. The reduction in the number of variables and constraints is 14% and 18% for the 2BB-1F approach and 26% and 35% for the 1BB-1F approach. It is important to highlight that the reduction is identical for all instances since the instance is an enlarged version of the same case study. Still, the connection between assets remains unchanged in all instances. It is also worth noting that the objective function is the same for every instance, regardless of the applied approach, indicating that they all represent the same energy problem. Finally, we cover a wide range of model sizes, from smaller instances representing optimisation models with a few thousand variables and constraints to larger instances with millions of variables and constraints.
@@ 5.2 Statistical analysis
The randomness in the results is due to the random number seed in the solver, which leads the solver to take different solution paths. Therefore, we can assume for the t-test that the results follow a normal probability distribution, the variances of the results are equal, and each individual in the population has an equal probability of being selected in the sample.
@@
The results presented in Table 3 indicate that 1BB-1F has a significantly different mean value than 2BB-2F in all instances. Additionally, Figure 7 shows that the average speedup values are higher for 1BB-1F in all cases. This implies that while 2BB-2F could potentially be faster than 1BB-1F in a particular instance, there is statistical evidence that, on average, 1BB-1F is faster. As for 2BB-1F, no statistical evidence suggests that it is faster than 2BB-F for half of the instances. However, it is worth noting that the larger case studies in the test, instances 5 and 6, show statistical evidence that the 2BB-1F approach is faster, on average, than the 2BB-2F.
@@ 5.3 Discussion
Nowadays, energy system modellers have access to several energy models that can be used to develop case studies. These models can have one or more available modelling options that can be framed in different approaches, which are described in this paper. Table 5 provides a general overview of the main available modelling approaches in a sample of energy system models. Models focusing on power systems mainly consist of two building blocks, nodes and energy assets. In contrast, most recent models that focus on multi-sector analysis, like Calliope and SpineOpt, have three building blocks: nodes, energy assets, and connections. The reason behind having more building blocks is to create a flexible model that can adapt to different energy asset configurations and represent different sectors. However, a flexible model can result in an increase in the number of variables and constraints, which can come with a computation cost. The results in the illustrative example in this paper estimate the impact in terms of speedup. As a general recommendation for energy system modellers, it is better to have fewer variables, especially for large-scale case studies.
@@
In order to reduce the size of the model, instead of reformulating the model using a single building block as proposed in this paper, some models provide the option to select specific methods tailored to the application. These methods also aim to reduce the number of variables and constraints. For instance, IRENA FlexTool gives users a selection of conversion and transfer methods, and some of those methods allow the use of fewer variables and constraints when the model is created, which also improves the solution process in the solver. FlexTool can actually present the example problem in 1BB-1F format with one extra variable using available methods and in full 1BB-1F format using carefully formulated user constraints, which can be entered as data. In both cases, the benefit comes from choosing a modelling approach or method with fewer variables and constraints to represent the same energy system. The advantage and strength of the 1BB-1F proposal in this paper is that it is a generic way to formulate the problem without depending on the definition of tailor-made methods that drops variables and constraints that are not used depending on the user settings in a given case study.
@@
The proposal for using only one building block with one flow variable (1BB-1F) is a new way of utilizing the network graph structure of energy systems. While this approach has been found to offer several advantages, it also poses two primary challenges and limitations. First and foremost, it involves an extra layer of abstraction as the proposal can establish direct connections between energy assets, which may not be very intuitive compared to traditional methods that use more building blocks like nodes and connections. However, this level of abstraction does not hinder the modellers from representing assets that perform the functions of nodes or connections, as evidenced in the case study where the same model, TulipaEnergyModel, was utilized to represent all the modelling approaches. Secondly, it is not always possible to obtain the benefits of asset-to-asset connections (1BB-1F) in situations where gas pressure constraints, transport delays, or DCOPF losses for transmission or distribution networks are considered. Modelling these situations results in variations between an energy asset’s incoming and outgoing flows, which reduces or eliminates the potential reduction on the model size of the 1BB-1F approach. In such cases, it becomes necessary to include two or four flow variables to represent the situation accurately. So, the 1BB-1F is not a silver bullet, but it can help in large-scale optimisation problems that can be simplified.
@@
| COMPETES [22] | Power Systems | 2004 | 2BB-2F |
| FlexTool [8] | Power Systems | 2021 | 2BB-1F |
@@
| PyPSA [13] | Multi-sector | 2018 | 1BB-1F |
@@ 6 Conclusion
Representing the same energy system in optimisation models with fewer variables and constraints without compromising accuracy is possible. Modellers must understand the impact of their choices on the model’s performance since flexible models allow for different approaches. This research has debunked the idea that an LP model’s simplicity cannot be improved without sacrificing fidelity and highlighted the importance of carefully considering different approaches representing energy system interactions. The extended concept of a network graph has shown great potential in reducing the number of variables and constraints in the model by using only one building block (1BB-1F). It can speed up the time required to create and solve the model by an average of 1.76 and 1.27, respectively, across all the case study instances. Notably, these benefits increase with the size of the problem, making it particularly advantageous for large-scale problems.
An interesting conclusion we found is that even though the solver can handle redundant and unnecessary constraints during the presolve procedures, providing it with fewer variables and constraints significantly improves solving time. This emphasizes the importance of making informed modelling choices.
We hope this analysis will raise awareness in the modelling community about the importance of making conscious decisions when representing energy systems. This analysis could be expanded to cover a wider range of physical dynamics of energy systems and their synergies with other commonly used reduction techniques in the literature.
@@ 9 Supplementary material
https://github.com/datejada/experiments-flexible-connection

### v3

@@ 3.2.2 Two building blocks with two flow variables (2BB-2F)
3.2.3 Two building blocks with one flow variable (2BB-1F)
This approach uses the energy assets and the balance nodes as building blocks. In addition, it uses a single free variable to represent the flow between nodes, which can take positive and negative values instead of two positive variables as in the previous approach. The main advantage of this method is that it eliminates the possibility of bidirectional flow between two nodes, as there is only one variable. However, it is not possible to model transmission or distribution network losses using a single free variable for the flow. Lastly, the free variable must have bounds in both directions to represent the capacity limits between the two nodes.
@@ 3.2.3 Two building blocks with one flow variable (2BB-1F)
5.
Non-negative variables:
3.2.4 One building block with one flow variable (1BB-1F)
This approach only uses the energy assets as building blocks. Using the graph-theory principles establishes the connection between energy assets as vertices and energy flows as edges. Connecting assets directly to each other (without any intervening nodes) can significantly reduce the number of variables and constraints required to represent the system. When transmission or distribution network losses are significant, two conversion assets can represent them—one for each direction with input-output ratios. A similar approach can be adopted for gas pressure flows. Therefore, this method enables us to easily model simple and complex situations, reducing the model size while retaining the same accuracy.
Figure 1(d) shows the battery with a renewable example for the 1BB-1F approach. Since the assets can connect among them, the battery can directly charge from renewable, and there is no need for an extra constraint to avoid charging from the grid. Therefore, we need 4 variables (3 flow variables + 1 storage level) and 6 constraints (4 capacity limits + 1 demand balance + 1 storage balance).
@@ 3.2.4 One building block with one flow variable (1BB-1F)
: flow from the battery to the electricity demand balance limited by battery capacity
: flow from the solar pv to the battery limited by battery capacity and the solar PV availability
: flow from the solar pv to the electricity demand balance limited by the solar availability
3.2.5 Summary
Table 1 summarises the number of variables and constraints per time step for each modelling approach in the example shown in Figure 1. It also shows the reduction in variables and constraints, with the 3BB-4F approach as a reference.
@@ 3.2.5 Summary
| Modelling Approach | Variables | Constraints | Non-zeros |
@@ 4.1 Energy system optimisation model
TulipaEnergyModel.jl is an optimisation model using 1BB-1F and determines the optimal investment and operation decisions for different types of assets (e.g., producers, consumers, conversion, storage, and transport). It is developed in Julia [16] and depends mainly on the JuMP.jl [17] and Graphs.jl [18] packages. The complete description of the model, its core concepts, mathematical formulation, and tutorials are available in the GitHub documentation of the model 333https://tulipaenergy.github.io/TulipaEnergyModel.jl/stable/.
4.2 Case study description
@@
Figure 2 shows an illustrative integrated energy system with three interconnected areas: Asgard, Midgard, and Valhalla. The diagram uses the 2BB-2F approach. In addition, it includes the flow and balance of electricity, heat, and gas within a mock-up energy grid to explore different possibilities for flows among energy assets using nodes. Asgard includes a combined cycle gas turbine, a solar photovoltaic installation, and a battery system. Midgard features a wind park, a hydro plant, and a small modular reactor for nuclear power generation. Valhalla focuses on hydrogen as an energy vector, with a hydrogen generator, a hydrogen storage facility, and a fuel cell. Transmission lines and gas pipelines connect the system, allowing energy transfer between areas.
Figures 3 and 4 show the equivalent energy system for the other two approaches. In general, each arrow represents a variable, and each BB element represents a constraint. Note that the number of flows (arrows) reduces compared to the 2BB-2F approach.
To assess the impact of varying problem sizes, we developed six instances of the problem, classified from small to large-scale optimisation. The smallest instance is labelled as 1, while the largest is labelled as 6. Each instance consists of hourly time steps, with differences in the time horizon covered. For example, instance 1 spans one month, whereas instance 6 covers four years.
Section 5 analyses the impact of these reductions in the model for each approach and instance. Finally, Section 10 has the link to the repository with the input data files for each approach.
@@ 5 Results
The results in this section were obtained using TulipaEnergyModel.jl version 0.6.1 and Gurobi version 11.0.0 on a 12th Gen Intel(R) Core(TM) i7-1255U 1.70 GHz processor and 16.0 GB RAM. Section 10 includes the link to the repository with the files to reproduce the experiments in this paper.
Table 2 shows the objective functions and problem size for each instance, where 2BB-2F is used as a reference in the table.
We cover a wide range of model sizes, from smaller instances representing optimisation models with a few thousand variables and constraints to larger instances with millions of variables and constraints.
The reduction in the number of variables and constraints is 14% and 18% for the 2BB-1F approach and 26% and 35% for the 1BB-1F approach. It is important to highlight that the reduction is identical for all instances since the instance is an enlarged version of the same case study. Still, the connection between assets remains unchanged in all instances. It is also worth noting that the objective function is the same for every instance, regardless of the applied approach, indicating that they all represent the same energy problem.
These results demonstrate that reducing the LP size while maintaining fidelity is possible. This finding is significant for practitioners looking to solve large-scale problems more efficiently without losing the desired level of detail.
@@ 5.2 Statistical analysis
The variability in the results stems from the random number seed used by the solver, which leads the solver to take different solution paths. Therefore, we can assume for the t-test that the results follow a normal probability distribution, the variances of the results are equal, and each individual in the population has an equal probability of being selected in the sample.
@@
The results presented in Table 3 indicate that 1BB-1F has a significantly different mean value than 2BB-2F in all instances. Additionally, Figure 7 shows that the average speedup values are higher for 1BB-1F in all cases. This implies that while 2BB-2F could potentially be faster than 1BB-1F in a particular instance, there is statistical evidence that, on average, 1BB-1F is faster. As for 2BB-1F, no statistical evidence suggests it is faster than 2BB-F for half of the instances. However, it is worth noting that the larger case studies in the test, instances 5 and 6, show statistical evidence that the 2BB-1F approach is faster, on average, than the 2BB-2F.
@@ 5.3 Discussion
Nowadays, energy system modellers have access to several energy models that can be used to develop case studies. Laveneziana et al. [20] present a critical review of energy planning models, comparing their capabilities and performance. However, this review does not explore the available modelling options described in this paper. Therefore, Table 5 provides a general overview of the main available modelling approaches in a sample of these energy system models. Models focusing on power systems mainly consist of two building blocks, nodes and energy assets. In contrast, most recent models that focus on multi-sector analysis, like Calliope and SpineOpt, have three building blocks: nodes, energy assets, and connections. The reason behind having more building blocks is to create a flexible model that can adapt to different energy asset configurations and represent different sectors. However, a flexible model can increase the number of variables and constraints, which can come with a computation cost. The results in the illustrative example in this paper estimate the impact in terms of speedup. As a general recommendation for energy system modellers, it is better to have fewer variables, especially for large-scale case studies.
@@
In order to reduce the size of the model, instead of reformulating the model using a single building block as shown in this paper, some models provide the option to select specific methods tailored to the application. These methods also aim to reduce the number of variables and constraints. For instance, IRENA FlexTool gives users a selection of conversion and transfer methods, and some of those methods allow the use of fewer variables and constraints when the model is created, which also improves the solution process in the solver. FlexTool can represent the example problem in 1BB-1F format with one extra variable using available methods and in full 1BB-1F format using carefully formulated user constraints, which can be entered as data.
Another example comes from PyPSA, a 2BB-1F approach model. In PyPSA, the bus serves as the fundamental building block, to which all energy assets—such as loads, generators, storage units, lines, transformers, and links—are connected. This design means that a direct connection between two energy assets is not possible, which distinguishes it from the 1BB-1F model. However, PyPSA provides the option to create custom constraints, allowing for more complex configurations, like those illustrated in the case study, while reducing the total number of constraints required. In both cases, Flextool and PyPSA, the benefit comes from choosing a modelling approach, method, or custom constraint that allows fewer variables and constraints to represent the same energy system. The advantage and strength of the proposed 1BB-1F in this paper is that it is a generic way to formulate the problem without depending on the definition of tailor-made methods or constraints that reduce model size depending on the user settings in a given case study.
@@
The proposal for using only one building block with one flow variable (1BB-1F) is a new way of utilising the network graph structure of energy systems. While this approach has been found to offer several advantages, it also poses two primary challenges and limitations. First and foremost, it involves an extra layer of abstraction as the proposal can establish direct connections between energy assets, which may not be intuitive compared to traditional methods that use more building blocks like nodes and connections. However, this level of abstraction does not hinder the modellers from representing assets that perform the functions of nodes or connections, as evidenced in the case study where the same model, where we used TulipaEnergyModel to represent all the modelling approaches. Secondly, it is not always possible to obtain all the benefits of asset-to-asset connections (1BB-1F) in situations where constraints to model gas pressure, heat transfer, or DCOPF with losses for transmission or distribution networks are considered. Modelling these situations results in variations between an energy asset’s incoming and outgoing flows, which partly reduces the potential savings on the model size of the 1BB-1F approach. In such cases, it becomes necessary to include two or four flow variables anyway to represent the situation accurately. Although the proposed 1BB-1F approach allows a direct connection between assets eliminating unnecessary elements in between, with their corresponding variables and constraints , it needs more variables to correctly represent more elaborated flows. Nevertheless, 1BB-1F helps large-scale optimisation problems that can be simplified.
@@
| COMPETES-TNO [23] | Power Systems | 2004 | 2BB-2F |
| FlexTool [8] | Power Systems | 2021 | 2BB-1F 444Models with extra methods or custom constraints that allow for reducing the number of variables and constraints, depending on the modelling needs. |
@@
| PyPSA [13] | Multi-sector | 2018 | 2BB-1F ††footnotemark: |
@@ 6 Conclusion
This paper demonstrates that large-scale LP energy system models can be reformulated to reduce computational burden without compromising fidelity. We introduce a graph-based formulation using a single building block—the energy asset—and show that this approach leads to a more compact and computationally efficient LP formulation. Compared with four traditional approaches that rely on multiple building blocks and connections, our method reduces variables and constraints by 26% and 35%, respectively, and achieves an average speedup of 1.27× in solving time. These improvements scale with problem size, offering substantial benefits for large-scale, high-resolution studies.
Our findings challenge the widespread belief that LP models cannot be improved without a loss of accuracy. Even though solvers can eliminate some redundancy during presolve, we show that providing a compact, well-formulated model leads to faster building and solving times. This is particularly valuable for researchers and practitioners working on regional or multi-sector energy system models.
Beyond technical benefits, this work advocates for more conscious modelling practices. Flexible models should not only accommodate various configurations but do so efficiently. The proposed approach enables scalable, high-fidelity modelling using existing solvers (software) and computers (hardware), making it easier to explore future energy pathways with greater resolution and technical details. We hope this research encourages the energy modelling community to reconsider assumptions about LP formulations and adopt modelling strategies that balance flexibility, fidelity, and performance.
@@ 7 CRediT author statement
8 Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used Grammarly to improve readability and language. After using this tool/service, the authors reviewed and edited the content as needed and took full responsibility for the publication’s content.
@@ 9 Acknowledgments
We would like to express our gratitude to the three anonymous reviewers for their valuable feedback, which significantly enhanced the initial version of this manuscript. Additionally, we extend our thanks to Fabian Neumann for his assistance in better understanding PyPSA’s formulation and capabilities, which were essential for the comments on PyPSA in this paper.
@@ 10 Supplementary material
The formulation presented in this paper is a simplified version of the TulipaEnergyModel.jl. The complete formulation, including all features of the model, can be found at the following link:
Annex A. Extended formulation
Section 3.1 shows a basic formulation for the proposed approach. In this section, we illustrate how to include other modelling features; for instance, here we show the equations to model the DC power flow and unit commitment problem.
A.1. DC power flow
For a flow connecting assets and , which belongs to the set , the power flow constraints are as follows:
Where:
: subset of flows that have DC power flow constraints
: voltage angle of asset in timestep . Notice that this asset is playing a role of a bus in the traditional energy modelling approach.
: per unit reactance between two assets, and
: base power, e.g., 100 MVA
A.2. Unit commitment
Assets within the subset will contain the unit commitment constraints in the model. The following constraints show a basic formulation using the unit commitment variable for the units on. However, more complex constraints with startup and shutdown variables, e.g., [27], can also be added to the model.
Where:
: subset of assets with unit commitment
: flow above the minimum operating point of asset in timestep
: flow variable from asset to asset in timestep
: units on variable of in timestep
: investment unit variable of asset
: initial number of units parameter of the asset
: maximum capacity parameter of the asset
: minimum capacity parameter of the asset